OpenStudy (anonymous):

simplify. sqrt of 315x^5. okay I got the next step... 3sqrt35x^5. what do I do with the power in the square root?

OpenStudy (anonymous):

\[3\sqrt{35x ^{5}}\]

OpenStudy (anonymous):

You divide the power by 2, because you are effectively raising the the initial 315x^5 to the power of 1/2. So it becomes 3 * sqrt(35) * x ^ (5/2)

OpenStudy (anonymous):

\[3\sqrt{35}x ^{5/2}\]

OpenStudy (anonymous):

nah just want to know what to do with the power in the root really.

OpenStudy (anonymous):

\[(315x ^{5})^{1/2}\]

OpenStudy (anonymous):

\[\sqrt{315}*(x ^{5})^{1/2}\]

OpenStudy (anonymous):

\[\sqrt{315x ^{5}} = \sqrt{315}*\sqrt{x ^{5}}= \sqrt{3*3*5*7}*\sqrt{x ^{2}*^{2}*x}\]

OpenStudy (anonymous):

\[3\sqrt{35}*x ^{2}\sqrt{x}\]

OpenStudy (anonymous):

\[3x ^{2}\sqrt{35x}\]

OpenStudy (anonymous):

ah, yes, 5 times a half... \[3\sqrt{35}*x ^{5/2}\] Ah!!! (again!) halfway... how did you go from \[\sqrt{x ^{2}*x ^{2}*x}\] to

OpenStudy (anonymous):

oops lol... \[x ^{2}\sqrt{x}\]

OpenStudy (anonymous):

when u have a square number under the root sign u take it out and write it once because root of a square is the number \[\sqrt{4}=\sqrt{2^{2}}=2\]

OpenStudy (anonymous):

Ah okay! Perfect! Thanks all!

OpenStudy (anonymous):

If u r satisfied with the answer, pls click GOOD ANSWER button for those u feel helped u....☺